LSP Co.’s stock price is $58.88, and it recently paid a $2.00 dividend. This dividend is expected to grow by 52% for the next 3 years, then grow forever at a constant rate, g; and rs = 12%. At what constant rate is the stock expected to grow after Year 3?


1
Expert's answer
2021-02-02T04:48:25-0500

Answer:


Step 1

Price of stock today is present value of dividend and intrinsic value of stock.


Step 2

Dividend today =2.0

Growth rate =52%

Dividend next =2(1.52)=3.04

D2=3.04(1.52)=4.62

D3=4.62(1.52)=7.02

D4=7.02(1+g)

P3=7.02(1+g)

P3= "7.02(1+g) \over (0.12-g)"

Value of stock today = "3.04 \over (1.12)"+ "4.62 \over (1.12)^2" + "7.02 \over (1.12)^3" + "7.02(1+g) \over (0.12-g)(1.12)^3"

 

 Step 3

58.88=2.71+3.68+4.99+"4.99(1+g) \over (0.12-g)"

47.5= "4.99(1+g) \over (0.12-g)"

9.51(0.12-g)=(1+g)

1.14-9.51g=1+g

0.14=10.51 g

g= "0.14 \over 10.51"

g=0.0134

g=1.34%


Therefore,1.34% is the constant rate that the stock expected to grow after Year 3.